# XATU: A Fine-grained Instruction-based Benchmark for Explainable Text Updates

### v1

@@ 4.2 Zero-shot Results
4.3 Fine-tuning Results
@@
4.4 Discussion
Figure 4: Boxplot comparing improvements of instruction fine-tuning vs. zero-shot performance.
@@ 4.4 Discussion
Figure 5: Boxplot comparing instruction-tuned model vs. pre-trained counterparts.
@@
As illustrated in the left side of Figure 4, we compare the results of three models after instruction fine-tuning with their zero-shot performance. We observe that the fine-tuned models consistently outperform their zero-shot counterparts, indicating the effectiveness of instruction fine-tuning in improving text editing performance. Across both coarse-grained and fine-grained instruction settings, we find that the Flan-UL2 model achieves the largest average improvements. On the other hand, Alpaca demonstrates superior robustness across all text editing tasks compared to Flan-T5 and Flan-UL2. This enhanced robustness may be attributed to the underlying decoder-only architecture.
@@
As depicted in the right side of Figure 4, the Flan-UL2 model consistently outperforms the UL2 model across all settings. Similarly, the Flan-T5 model performs better than the base T5 model. These results highlight the effectiveness of instruction tuning in improving the performance of language models across different settings. The instruction-tuning procedure allows the models to better understand and follow the provided instructions, resulting in more accurate and appropriate text edits. We also notice edits generated with coarse-grained instructions show better robustness across all tasks.
5 Conclusion
@@
This paper introduces XATU, the first benchmark for explainable text updates with fine-grained instructions. XATU is a diverse benchmark covering a wide range of topics and text types and leverages high-quality data sources from various existing sources. We compare existing open and closed instruction-tuned language models under both the zero-shot and fine-tuning settings and real their capabilities to edit text and follow instructions. By releasing the benchmark to the community, we hope to stimulate further research in developing instruction-based text editing models and even potentially interactive text generation systems.
Limitations
Our benchmark XATU still has several limitations. First, XATU only includes data sources from the English language. We leave evaluating multilingual text editing capabilities to future work.
Second, XATU only has thousands of examples due to budget limitations. However, experiment results demonstrate that instruction fine-tuning with $200$ high-quality examples already shows significant improvements.
Third, we only compare LLMs performance under zero-shot and fine-tuning settings. We leave in-context learning results (few-shot) to future work.
@@ Appendix A Data Annotation
{tcolorbox}
[fontupper=, title=Prompts to generate fine-grained instruction] [no prose]
# Task:
Your task is to write a detailed instruction that enables the AI assistant to edit the original text into a revised text based on the references. The instruction must not cause information leakage about the revised text.
# Original text:
{{input}}
# Reference:
{{reference}}
# Revised text:
{{output}}
# Instruction:
{tcolorbox}
[fontupper=, title=Prompts to generate explanation] [no prose]
# Task:
Your task is to provide a two-sentence explanation of the edits made based on the instruction and reference by comparing the original and revised texts.
# Instruction:
{{instruction}}
# Original text:
{{input}}
# Reference:
{{reference}}
# Revised text:
{{output}}
# Explanation:
@@ Appendix B Implementation Details
{tcolorbox}
[fontupper=, title=Prompts for Text Editing] [no prose]
Below is an instruction that describes a task, along with an input text paired with a reference and an explanation that provide further context. Please edit the input text based on the instruction, the reference, and the explanation. Your response should only include the edited output.
# Instruction:
{{instruction}}
# Input:
{{input}}
# Reference:
{{reference}}
# Explanation:
{{explanation}}
# Response:

### v2

@@ 4.2.   Zero-shot Results
4.3.   Fine-tuning Results
@@
4.4.   Discussion
Figure 4: Fine-tuning with fine-grained instructions (-fine) vs. coarse instructions (-c).
@@ 4.4.   Discussion
Figure 5: Boxplot comparing instruction-tuned LLMs (Flan-xx) vs. pre-trained counterparts with fine-grained (-fine) and coarse instructions (-c).
@@
As illustrated in Figure 4, we compare the results of three models after instruction fine-tuning with fine-grained instructions in XATU vs. coarse-grained instructions. We observe that fine-grained instruction models consistently outperform their coarse-grained counterparts, indicating the effectiveness of detailed instruction in improving instruction following and text editing capabilities. Across both coarse-grained and fine-grained instruction settings, we find that the Flan-UL2 model achieves the largest average improvements. On the other hand, Alpaca demonstrates superior robustness across all text editing tasks compared to Flan-T5 and Flan-UL2.
@@
As depicted in Figure 5, the Flan-UL2 model consistently outperforms the UL2 model across all settings. Similarly, the Flan-T5 model performs better than the base T5 model. These results highlight the effectiveness of instruction tuning in improving the performance of language models across different settings. The instruction-tuning procedure allows the models to better understand and follow the provided instructions, resulting in more accurate and appropriate text edits. We also notice edits generated with coarse-grained instructions show better robustness across all tasks.
5.   Conclusion
@@
This paper introduces XATU, the first benchmark for explainable text updates with fine-grained instructions. XATU is a diverse benchmark covering a wide range of topics and text types and leverages high-quality data sources from various existing sources. We compare existing open and closed instruction-tuned language models under both the zero-shot and fine-tuning settings and reveal their capabilities to edit text and follow instructions. By releasing the benchmark to the community, we hope to stimulate further research in developing instruction-based text editing models and even potentially interactive text generation systems.
6.   Acknowledgements
We thank the anonymous reviewers for their valuable reviews. Additionally, we appreciate Tom Mitchell and the entire team at Megagon Labs for insightful discussions.
7.   Bibliographical References
[…]
  • Akyürek et al. (2023) Afra Feyza Akyürek, Ekin Akyürek, Aman Madaan, Ashwin Kalyan, Peter Clark, Derry Wijaya, and Niket Tandon. 2023. Rl4f: Generating natural language feedback with reinforcement learning for repairing model outputs. arXiv preprint arXiv:2305.08844.
  • Alva-Manchego et al. (2020) Fernando Alva-Manchego, Louis Martin, Antoine Bordes, Carolina Scarton, Benoît Sagot, and Lucia Specia. 2020. ASSET: A dataset for tuning and evaluation of sentence simplification models with multiple rewriting transformations. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4668–4679, Online. Association for Computational Linguistics.
  • Anthonio et al. (2020) Talita Anthonio, Irshad Bhat, and Michael Roth. 2020. wikiHowToImprove: A resource and analyses on edits in instructional texts. In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 5721–5729, Marseille, France. European Language Resources Association.
  • Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862.
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Advances in Neural Information Processing Systems, volume 33, pages 1877–1901. Curran Associates, Inc.
  • Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Eric Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, et al. 2022. Scaling instruction-finetuned language models. arXiv preprint arXiv:2210.11416.
  • DiMarco and Hirst (1993) Chrysanne DiMarco and Graeme Hirst. 1993. A computational theory of goal-directed style in syntax. Computational Linguistics, 19(3):451–500.
  • Du et al. (2022) Wanyu Du, Vipul Raheja, Dhruv Kumar, Zae Myung Kim, Melissa Lopez, and Dongyeop Kang. 2022. Understanding iterative revision from human-written text. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 3573–3590, Dublin, Ireland. Association for Computational Linguistics.
  • Dwivedi-Yu et al. (2022) Jane Dwivedi-Yu, Timo Schick, Zhengbao Jiang, Maria Lomeli, Patrick Lewis, Gautier Izacard, Edouard Grave, Sebastian Riedel, and Fabio Petroni. 2022. Editeval: An instruction-based benchmark for text improvements. arXiv preprint arXiv:2209.13331.
  • Faltings et al. (2021) Felix Faltings, Michel Galley, Gerold Hintz, Chris Brockett, Chris Quirk, Jianfeng Gao, and Bill Dolan. 2021. Text editing by command. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 5259–5274, Online. Association for Computational Linguistics.
  • Faruqui et al. (2018) Manaal Faruqui, Ellie Pavlick, Ian Tenney, and Dipanjan Das. 2018. WikiAtomicEdits: A multilingual corpus of Wikipedia edits for modeling language and discourse. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 305–315, Brussels, Belgium. Association for Computational Linguistics.
  • Guu et al. (2018) Kelvin Guu, Tatsunori B. Hashimoto, Yonatan Oren, and Percy Liang. 2018. Generating sentences by editing prototypes. Transactions of the Association for Computational Linguistics, 6:437–450.
  • Honovich et al. (2023) Or Honovich, Uri Shaham, Samuel R. Bowman, and Omer Levy. 2023. Instruction induction: From few examples to natural language task descriptions. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1935–1952, Toronto, Canada. Association for Computational Linguistics.
  • Hsieh et al. (2023) Cheng-Yu Hsieh, Chun-Liang Li, Chih-Kuan Yeh, Hootan Nakhost, Yasuhisa Fujii, Alexander Ratner, Ranjay Krishna, Chen-Yu Lee, and Tomas Pfister. 2023. Distilling step-by-step! outperforming larger language models with less training data and smaller model sizes. arXiv preprint arXiv:2305.02301.
  • Hu et al. (2022) Edward J Hu, yelong shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-rank adaptation of large language models. In International Conference on Learning Representations.
  • Iso et al. (2020) Hayate Iso, Chao Qiao, and Hang Li. 2020. Fact-based Text Editing. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 171–182, Online. Association for Computational Linguistics.
  • Iv et al. (2022) Robert Iv, Alexandre Passos, Sameer Singh, and Ming-Wei Chang. 2022. FRUIT: Faithfully reflecting updated information in text. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 3670–3686, Seattle, United States. Association for Computational Linguistics.
  • Li et al. (2023) Zizhong Li, Haopeng Zhang, and Jiawei Zhang. 2023. A revisit of fake news dataset with augmented fact-checking by chatgpt. arXiv preprint arXiv:2312.11870.
  • Liu et al. (2022) Yixin Liu, Budhaditya Deb, Milagro Teruel, Aaron Halfaker, Dragomir Radev, and Ahmed H Awadallah. 2022. On improving summarization factual consistency from natural language feedback. arXiv preprint arXiv:2212.09968.
  • Lyu et al. (2021) Yiwei Lyu, Paul Pu Liang, Hai Pham, Eduard Hovy, Barnabás Póczos, Ruslan Salakhutdinov, and Louis-Philippe Morency. 2021. StylePTB: A compositional benchmark for fine-grained controllable text style transfer. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 2116–2138, Online. Association for Computational Linguistics.
  • Malmi et al. (2019) Eric Malmi, Sebastian Krause, Sascha Rothe, Daniil Mirylenka, and Aliaksei Severyn. 2019. Encode, tag, realize: High-precision text editing. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 5054–5065, Hong Kong, China. Association for Computational Linguistics.
  • Mukherjee et al. (2023) Subhabrata Mukherjee, Arindam Mitra, Ganesh Jawahar, Sahaj Agarwal, Hamid Palangi, and Ahmed Awadallah. 2023. Orca: Progressive learning from complex explanation traces of gpt-4. arXiv preprint arXiv:2306.02707.
  • Napoles et al. (2017) Courtney Napoles, Keisuke Sakaguchi, and Joel Tetreault. 2017. JFLEG: A fluency corpus and benchmark for grammatical error correction. In Proceedings of the 15th Conference of the European Chapter of the Association for Computational Linguistics: Volume 2, Short Papers, pages 229–234, Valencia, Spain. Association for Computational Linguistics.
  • Narayan et al. (2018) Shashi Narayan, Shay B. Cohen, and Mirella Lapata. 2018. Don’t give me the details, just the summary! topic-aware convolutional neural networks for extreme summarization. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 1797–1807, Brussels, Belgium. Association for Computational Linguistics.
  • OpenAI (2023) OpenAI. 2023. Gpt-4 technical report.
  • Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. arXiv preprint arXiv:2203.02155.
  • Pryzant et al. (2020) Reid Pryzant, Richard Diehl Martinez, Nathan Dass, Sadao Kurohashi, Dan Jurafsky, and Diyi Yang. 2020. Automatically neutralizing subjective bias in text. In Proceedings of the aaai conference on artificial intelligence, volume 34, pages 480–489.
  • Radford et al. (2019) Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. 2019. Language models are unsupervised multitask learners. Technical report, Open AI.
  • Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, Peter J Liu, et al. 2020. Exploring the limits of transfer learning with a unified text-to-text transformer. J. Mach. Learn. Res., 21(140):1–67.
  • Reid and Neubig (2022) Machel Reid and Graham Neubig. 2022. Learning to model editing processes. In Findings of the Association for Computational Linguistics: EMNLP 2022, pages 3822–3832, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
  • Schick et al. (2023) Timo Schick, Jane A. Yu, Zhengbao Jiang, Fabio Petroni, Patrick Lewis, Gautier Izacard, Qingfei You, Christoforos Nalmpantis, Edouard Grave, and Sebastian Riedel. 2023. PEER: A collaborative language model. In The Eleventh International Conference on Learning Representations.
  • Spangher et al. (2022) Alexander Spangher, Xiang Ren, Jonathan May, and Nanyun Peng. 2022. NewsEdits: A news article revision dataset and a novel document-level reasoning challenge. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 127–157, Seattle, United States. Association for Computational Linguistics.
  • Taori et al. (2023) Rohan Taori, Ishaan Gulrajani, Tianyi Zhang, Yann Dubois, Xuechen Li, Carlos Guestrin, Percy Liang, and Tatsunori B Hashimoto. 2023. Alpaca: A strong, replicable instruction-following model. Stanford Center for Research on Foundation Models. https://crfm. stanford. edu/2023/03/13/alpaca. html, 3(6):7.
  • Tay et al. (2023) Yi Tay, Mostafa Dehghani, Vinh Q. Tran, Xavier Garcia, Jason Wei, Xuezhi Wang, Hyung Won Chung, Dara Bahri, Tal Schuster, Steven Zheng, Denny Zhou, Neil Houlsby, and Donald Metzler. 2023. UL2: Unifying language learning paradigms. In The Eleventh International Conference on Learning Representations.
  • Thorne and Vlachos (2021) James Thorne and Andreas Vlachos. 2021. Evidence-based factual error correction. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 3298–3309, Online. Association for Computational Linguistics.
  • Thorne et al. (2018) James Thorne, Andreas Vlachos, Christos Christodoulopoulos, and Arpit Mittal. 2018. FEVER: a large-scale dataset for fact extraction and VERification. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), pages 809–819, New Orleans, Louisiana. Association for Computational Linguistics.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.
  • Veselovsky et al. (2023) Veniamin Veselovsky, Manoel Horta Ribeiro, and Robert West. 2023. Artificial artificial artificial intelligence: Crowd workers widely use large language models for text production tasks.
  • Wei et al. (2022) Jason Wei, Maarten Bosma, Vincent Zhao, Kelvin Guu, Adams Wei Yu, Brian Lester, Nan Du, Andrew M. Dai, and Quoc V Le. 2022. Finetuned language models are zero-shot learners. In International Conference on Learning Representations.
  • Welleck et al. (2023) Sean Welleck, Ximing Lu, Peter West, Faeze Brahman, Tianxiao Shen, Daniel Khashabi, and Yejin Choi. 2023. Generating sequences by learning to self-correct. In The Eleventh International Conference on Learning Representations.
  • Xu et al. (2016) Wei Xu, Courtney Napoles, Ellie Pavlick, Quanze Chen, and Chris Callison-Burch. 2016. Optimizing statistical machine translation for text simplification. Transactions of the Association for Computational Linguistics, 4:401–415.
  • Yang et al. (2017) Diyi Yang, Aaron Halfaker, Robert Kraut, and Eduard Hovy. 2017. Identifying semantic edit intentions from revisions in wikipedia. In Proceedings of the 2017 Conference on Empirical Methods in Natural Language Processing, pages 2000–2010.
  • Yin et al. (2019) Pengcheng Yin, Graham Neubig, Miltiadis Allamanis, Marc Brockschmidt, and Alexander L. Gaunt. 2019. Learning to represent edits. In International Conference on Learning Representations.
  • Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Richmond Y. Wong, Bjoern Hartmann, and Qian Yang. 2023. Why johnny can’t prompt: How non-ai experts try (and fail) to design llm prompts. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23, New York, NY, USA. Association for Computing Machinery.
  • Zhang et al. (2023a) Haopeng Zhang, Xiao Liu, and Jiawei Zhang. 2023a. Extractive summarization via ChatGPT for faithful summary generation. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 3270–3278, Singapore. Association for Computational Linguistics.
  • Zhang et al. (2023b) Haopeng Zhang, Xiao Liu, and Jiawei Zhang. 2023b. SummIt: Iterative text summarization via ChatGPT. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 10644–10657, Singapore. Association for Computational Linguistics.
  • Zhang et al. (2022) Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi Victoria Lin, Todor Mihaylov, Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit Singh Koura, Anjali Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022. Opt: Open pre-trained transformer language models.
  • Zhao et al. (2018) Sanqiang Zhao, Rui Meng, Daqing He, Andi Saptono, and Bambang Parmanto. 2018. Integrating transformer and paraphrase rules for sentence simplification. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 3164–3173, Brussels, Belgium. Association for Computational Linguistics.
  • Zhong (2021) Yang Zhong. 2021. WIKIBIAS: Detecting Multi-Span Subjective Biases in Language. Ph.D. thesis, The Ohio State University.
  • Zhou et al. (2023a) Chunting Zhou, Pengfei Liu, Puxin Xu, Srini Iyer, Jiao Sun, Yuning Mao, Xuezhe Ma, Avia Efrat, Ping Yu, Lili Yu, et al. 2023a. Lima: Less is more for alignment. arXiv preprint arXiv:2305.11206.
  • Zhou et al. (2023b) Yongchao Zhou, Andrei Ioan Muresanu, Ziwen Han, Keiran Paster, Silviu Pitis, Harris Chan, and Jimmy Ba. 2023b. Large language models are human-level prompt engineers. In The Eleventh International Conference on Learning Representations.
@@ Appendix A Data Annotation
Prompts to generate fine-grained instruction
[no prose] # Task: Your task is to write a detailed instruction that enables the AI assistant to edit the original text into a revised text based on the references. The instruction must not cause information leakage about the revised text. # Original text: {{input}} # Reference: {{reference}} # Revised text: {{output}} # Instruction:
Prompts to generate explanation
[no prose] # Task: Your task is to provide a two-sentence explanation of the edits made based on the instruction and reference by comparing the original and revised texts. # Instruction: {{instruction}} # Original text: {{input}} # Reference: {{reference}} # Revised text: {{output}} # Explanation:
@@ Appendix B Implementation Details
Prompts for Text Editing
[no prose] Below is an instruction that describes a task, along with an input text paired with a reference and an explanation that provides further context. Please edit the input text based on the instructions, the reference, and the explanation. Your response should only include the edited output. # Instruction: {{instruction}} # Input: {{input}} # Reference: {{reference}} # Explanation: {{explanation}} # Response: